OpenStudy (anonymous):

whats the answer Factor completely: 50x2 – 32 Answer (5x – 8)(10x + 4) (5x – 8)(10x – 4) 2(5x – 4)(5x – 4) 2(5x – 4)(5x + 4)

OpenStudy (anonymous):

Take out the common factor, then use the identity \(a^2-b^2 = (a+b)(a-b)\) to factorize it.

OpenStudy (anonymous):

i have no clue what any of that means can u just give me the answer

OpenStudy (anonymous):

Well, the common factor between 50 and 32 is 2, so take 2 out. You'd get this: 2(25x^2 - 16) You can break these perfect squares into the form: (a+b)(a-b) Therefore, your answer is 2(5x+4)(5x-4)

OpenStudy (anonymous):

thats not even a choice

OpenStudy (anonymous):

That's the same as your last choice: 2(5x-4)(5x+4) It's the same thing, only reversed. Just like how 2+6 = 8 and 6+2 = 8